Question 4: Inventory Valuation
Suggested time: 15 minutes
Athabasca Auto makes and sell auto batteries. In 2019 it made 120,000 batteries and sold 95,000 of them, at an average selling price of $480 per unit. The following additional information relates to Athabasca Auto for 2019:
Direct Materials
$160.00 per unit
Direct manufacturing labour
$32.00 per unit
Variable manufacturing costs
$64.00 per unit
Sales commissions
$48.00 per part
Fixed manufacturing costs
$1,750,000 per year
Fixed administrative expenses
$670,000 per year
What is Athabasca Auto manufacturing cost per unit using:
a. Variable costing (7 marks)
b. Absorption costing (8 marks)
Answer #1

Part A

Calculation of Manufacturing Cost per Unit as per Variable Costing

Manufacturing Cost per Unit = Direct Materials Cost per Unit + Direct Manufacturing Labour per Unit + Variable Manufacturing Cost per Unit

Direct Materials Cost per Unit = $ 160

Direct Manufacturing Labour per Unit = $ 32

Variable Manufacturing Cost per Unit = $ 64

Manufacturing Cost per Unit = 160 + 32 + 64

Manufacturing Cost per Unit = $ 256

Part B

Calculation of Manufacturing Cost per Unit as per Absorption Costing

Manufacturing Cost per Unit = Direct Materials Cost per Unit + Direct Manufacturing Labour per Unit + Variable Manufacturing Cost per Unit + Fixed Manufacturing Cost per Unit

Direct Materials Cost per Unit = $ 160

Direct Manufacturing Labour per Unit = $ 32

Variable Manufacturing Cost per Unit = $ 64

Fixed Manufacturing Costs Per Unit = Fixed Manufacturing Costs / 120,000 Units

Fixed Manufacturing Cost per Unit = 1,750,000 / 120,000 = $ 14.58 Per Unit

Manufacturing Cost per Unit = 160 + 32 + 64 + 14.58

Manufacturing Cost per Unit = $ 270.58